

********************************************************************
UARS WEEKLY STATUS REPORT - 23 JULY 1999
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM's operations this past week were normal, without apparent
problems. The instrument was turned off on July 22 as part of
power-sharing.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
The PEM/AXIS activated between June 30 and July 7, and again between
July 16 and July 21. While AXIS was active, it produced energy deposition
from 16 pixels based on precipitating electrons derived from measured X-ray
fluxes.
The PEM spectrometers continue to be active under PMON control. While
under PMON control, the PEM instrument computed energy deposition from its
particle spectrometers.
The FOT and the PEM team are completing close-out documentation on
the HEPS Energy Processor trouble; however progress has been hampered
with the departure of Peter Thomas.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters and
science channel data within expected operating ranges. Science data
products continue to be of excellent quality.
***********
MLS
***********
MLS was reactivated successfully on July 22 and is scheduled to operate through
July 27.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
o DAAC data created in the month of May 1999 has been previewed for
HALOE, PEM, NMC, and UKMO.
o The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
19 days CLA (7.0) PROD (L1-L3) (473R-125C-11F=337L)
9 days GPI (4.05) PROD (CLAES L3B) (43R-9C-1F=33L)
49 days WIN (5.11) PROD (L1-L3) (1540R-994C-17F=529L)
CDHF pending software upgrades; DEC AXP Fortran 6.5, DEC AXP OpenVMS 6.2,
OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 10 July 1999 (Orbit 42773, GMT Day 191) through 16 July
(GMT Day 197).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 195/0036 42833 On
WINDII 196/1147 42855 On/Off
HRDI 196/2122 42861 Low Power
PEM 197/1336 42871 AXIS On
The beta angle ranged between the angles of 5.1 degrees to 21.8 degrees for
this report period. The beta angle is now increasing toward a maximum
angle of 75 degrees on 31 July 1999 (DOY 212).
A Reverse-to-Forward Yaw Around maneuver (number 80) was successfully
executed in Orbit 42782 beginning 191/13:58:35 and completing at
191/14:35:45 on 10 July 1999. The SSPP Instruments (ACRIM, SOLSTICE,
SUSIM) were enabled as well as HRDI and PEM in day only operations. Due to
the SA position at 269 degrees, the effective night length of the
spacecraft after the maneuver was approximately 53 minutes. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night load
bus voltage was 23.8 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 24.5 and 24.7 percent respectively.
The next (forward-to-reverse) YAW is scheduled for 20 August 1999 (DOY
232).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.4 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.1 degrees C. Additional battery performance data
for each battery on 09 July 1999 and 16 July 1999 are provided in Appendix
D.
The clock error ranged from -4.5 to 0.9 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 08
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 01
% HALOE 01
% WINDII 01
% HRDI 01
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
Two (2) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/179 - 2287.496941
Transponder B: 99/180 - 2287.498977
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/196 show the observatory data loss to be
24 hours, 54 minutes, 18.068 seconds (an increase of 18.592 seconds since
the last report period). This is a 0.0364 percent data loss which equals a
99.9636 percent data capture for the mission.
The increase of 18.592 seconds occurred over three (3) day:
7 0.224 seconds on Day 190 due to a data hit,
7 8.800 seconds on Day 192 due to DRC not covered,
7 9.568 seconds on Day 194 due to a data hit and no overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The FOT received a set of Y2K planning aids from FDF which will be used to
facilitate UARS Y2K end-to-end testing. Eleven 3-day planning aid files
were received and processed by the FOT. Two of these files did not process
successful on the CMS due to either incorrect data, or invalid data
formats. FDF has been notified of the problems, and are in the process of
correcting the files.
The FOT participated in an NCC 99 Y2K interface test on 13 July 1999. The
purpose of the test was to simulate a year 1999-to-2000 rollover and to
test how the NCC 99 software would handle scheduling tasks in the new year.
Minor scheduling problems were experienced, but most seemed to be either
due to mis-configurations between the various test entities, or because of
problems with the "test" software version on the User Planning System.
MEETINGS
None
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 20
August 1999 (DOY 232).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 July 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42781 191/1330 99-238 OBC Yaw around
42833 195/0036 99-239 HALOE Power ON
42855 196/1147 99-241 WINDII Power ON/OFF due to low LBV
42861 196/2122 99-242 HRDI Power ON, Low pwr.
42871 197/1336 99-225 PEM AXIS Power On
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 16 July 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRPSEQBIMON Y-HI HRDI QUIET 42777-8,85
HATBLT3 Y-LO Tel Barrel temp 42834
HAH28VA R-LO HALO HTR/M 42834,5,8
HASMIRRT Y-HI HALO SECON 42835
PEPCUPRIIE R-HI PCU CURREN 42853
OBSVLBEON R-LO EON Load B 42856
OBSBATVEON Y-LO Battery 2 42856
OBSBATVEON Y-LO Battery 3 42856
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 16 July 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
None this report period
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 16 July 1999
09 July 1999 - GMT Day 190 Beta = 5.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -100.8/-123.2 -3.05/-4.31 +0.0/+0.0 16.5 0.000 0.0
2 +33.6/-22.4 +5.47/+4.51 +30.0/-13.6 24.6 1.015 21.4
3 +50.4/-5.6 +4.51/+3.55 +29.6/-14.4 24.5 1.016 21.1
16 July 1999 - GMT Day 197 Beta = 21.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -100.8/-123.2 -3.05/-4.31 +0.0/ +0.0 16.5 0.000 0.0
2 +44.8/-28.0 +5.79/+4.83 +30.4/-14.4 25.0 1.023 21.8
3 +28.0/-33.6 +4.83/+3.87 +29.6/-14.8 24.8 1.023 21.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 16 July 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-034 HRDI 42856 196/1323 N/A TMON1 Trip due to low power
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-033 FOT see list July 99 14170 Generic Late Acq.
99-034 HRDI 42856 196/1323 N/A TMON1 Trip due to low power
SEVNTFW99/29